A stone of mass $10$g placed at the top of a tower $50$m high is allowed to fall freely. Show that law of conservation of energy is holds.
1
Expert's answer
2020-06-17T09:39:13-0400

Given, Mass of stone is "M=10g" ,Height of the tower is "H=50m" .


Initially, energy on system is just form of potential in particularly gravitional potential energy,thus


"E=MgH"

Now, consider at any arbitrary time ,the stone reach to distance "X" and velocity is "v" ,thus kinetic energy at that moment is

"K=\frac{1}{2}Mv^2"

Potential energy is

"U=Mg(H-X)"

Also, due to free fall under gravity

"v^2=2gX"

Hence, considering the above fact energy on the system will be

"K+U\\
\implies K+ U=MgH=E"

We are done.